# CITYWALK: Enhancing LLM-Based C++ Unit Test Generation via Project-Dependency Awareness and Language-Specific Knowledge

### v1

@@ 4.2. Answering RQ2
4.3. Answering RQ3
To answer this question, We further employ CITYWALK with two open-source code LLMs, namely CodeGeeX4-9B and DeepSeek-V2.5-236B. Specifically, we conduct ablation experiments on each LLM individually to investigate the impact of querying the corresponding LLM using the prompting strategy and additional contextual guidance designed by CITYWALK. Table 10, Table 11, Table 12 and Table 13 present the comparison results using the correctness and coverage metrics, respectively. Each LLM’s results are displayed in two lines: the first line shows the results when the LLM directly utilizes the basic prompt to generate unit test cases for the given focal methods, and the second line presents the results when integrated with CITYWALK. We derive two key insights from the statistical results within the tables: Integrating CITYWALK as a complementary plug-in enhances the performance of C++ unit test generation. As observed, both code LLMs demonstrate consistent improvements in all four evaluation metrics across the eight projects. The performance gains exhibit a marked increase with the number of parameters. The average performance gain of each metric (listed in the Avg. $\uparrow$ column) with DeepSeek-V2.5-236B is substantially greater than that of CodeGeeX4-9B.
Table 10. Generalization Results of CITYWALK on Different Code LLMs in Terms of Compilation Success Rate (CSR)
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg. $\uparrow$
CodeGeeX4 0.00% 6.95% 38.98% 0.09% 4.22% 11.75% 10.70% 35.75% 21.96%
CodeGeeX4-9B w/ CITYWALK 43.59% 42.11% 54.64% 4.30% 31.45% 29.12% 36.45% 42.44%
DeepSeek-V2.5 0.00% 0.00% 17.79% 0.00% 7.35% 13.55% 25.14% 51.93% 40.39%
DeepSeek-V2.5-236B w/ CITYWALK 74.62% 72.67% 43.45% 47.83% 41.56% 42.53% 53.89% 62.34%
Table 11. Generalization Results of CITYWALK on Different Code LLMs in Terms of Execution Pass Rate (EPR)
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg. $\uparrow$
CodeGeeX4 0.00% 6.00% 2.76% 0.09% 3.82% 0.36% 8.96% 15.54% 22.84%
CodeGeeX4-9B w/ CITYWALK 30.77% 37.92% 42.92% 2.15% 26.78% 25.56% 21.78% 32.34%
DeepSeek-V2.5 0.00% 0.00% 13.78% 0.00% 6.69% 12.02% 24.02% 39.91% 35.46%
DeepSeek-V2.5-236B w/ CITYWALK 67.69% 64.92% 31.82% 42.61% 36.78% 37.63% 47.12% 51.56%
Table 12. Generalization Results of CITYWALK on Different Code LLMs in Terms of Line Coverage ($\mathbf{Cov_{L}}$)
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg. $\uparrow$
CodeGeeX4 0.00% 15.30% 16.54% 0.00% 8.44% 10.45% 7.14% 21.75% 17.75%
CodeGeeX4-9B w/ CITYWALK 15.78% 46.15% 35.28% 28.04% 19.12% 12.45% 26.89% 37.89%
DeepSeek-V2.5 0.00% 0.00% 31.50% 0.00% 20.94% 22.23% 37.31% 30.04% 26.47%
DeepSeek-V2.5-236B w/ CITYWALK 22.44% 78.17% 42.80% 39.71% 52.23% 26.63% 48.35% 43.45%
Table 13. Generalization Results of CITYWALK on Different Code LLMs in Terms of Branch Coverage ($\mathbf{Cov_{B}}$)
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg. $\uparrow$
CodeGeeX4 0.00% 9.97% 15.73% 0.00% 9.18% 7.88% 10.81% 22.22% 12.68%
CodeGeeX4-9B w/ CITYWALK 5.18% 33.07% 31.23% 21.09% 21.45% 10.72% 21.53% 32.98%
DeepSeek-V2.5 0.00% 0.00% 24.24% 0.00% 19.03% 23.24% 32.58% 28.89% 23.65%
DeepSeek-V2.5-236B w/ CITYWALK 23.65% 65.98% 35.06% 31.94% 51.89% 25.15% 45.41% 38.12%
Answer to RQ3: CITYWALK effectively harnesses the latent intelligence of the LLMs, demonstrating the potential for seamless integration with additional LLMs in a plug-and-play manner.
@@ 4.3. Answering RQ3
5. Discussion
5.1. Usefulness of CITYWALK in Finding Real-World Bugs
Finding real software bugs is one of the most important criteria for evaluating an automated unit test generation approach. Therefore, we conduct a case-by-case analysis to assess the ability of CITYWALK to trigger exceptions and discover bugs. First, we select two bugs from tinyxml2 that were identified by CITYWALK-generated test cases. These are real-world issues that have been reported by human developers. As illustrated in Figure 11(a), the CITYWALK-generated test case first creates an XMLDocument object, doc, which then invokes the Value function. This function, in turn, calls XMLNode::Value(). Since the variable _value within XMLNode::Value() may not have been initialized, it triggers an assertion error in the GetStr function. This bug, identified by CITYWALK, corresponds to Issue #323  [^12] https://github.com/leethomason/tinyxml2/issues/323 in tinyxml2, where a human developer also commits an assertion failure when calling Value. Figure 11(b) shows another bug triggered by CITYWALK, where invalid hexadecimal format string arguments are passed to ToInt64. This bug is associated with Issue #825  [^13] https://github.com/leethomason/tinyxml2/issues/825 in tinyxml2, where a human developer also identifies a similar issue when calling the static member function ToInt64. In summary, the test cases generated by CITYWALK are effective in identifying real bugs within open-source projects.
@@ 5.2. False-Positive Executable Test Cases
Existing LLM-based unit test generation approaches (Yuan et al., 2024; Wang et al., 2024b) primarily rely on compiler error messages to guide LLMs in an iterative fixing process after generated test cases fail to execute. CITYWALK excludes such a fixing strategy because the error messages inherently contain the actual outcomes of the focal methods. This can lead LLMs to adjust expected values based on actual results or manipulate outputs to match expected outcomes, thereby passing tests under misleading circumstances. As illustrated in Figure 12, when provided with the error message, GPT-4o directly uses the expected value "str" as the input for assertions, resulting in a passing test case that is, in reality, a false positive. This behavior is undesirable, as it undermines the reliability of unit testing and fails to verify the intended functionality of the focal method. Thus, effectively addressing the false positive issue remains as our future work.
@@ 5.3. Threats to Validity
The primary threat to the validity of CITYWALK lies in the diversity of the projects used for evaluation and its generalization to other LLMs. In this paper, we collect 1057 focal methods across eight real-world open-source C++ projects crawled from GitHub, ensuring a degree of diversity and quality in our evaluation. For comparison, we select two open-source code LLMs (CodeGeeX4 and DeepSeek-V2.5) and two closed-source commercial LLMs (ChatGPT and GPT-4o), considering their varied model sizes, architectures, and effectiveness on coding tasks. Future work will involve expanding the benchmark and integrating additional LLMs to better assess the generalizability of CITYWALK.
@@ 6.3. LLM-based Automated Unit Test Generation
With the emergence of generative artificial intelligence, researchers have increasingly investigated LLM-driven approaches for automatically generating unit test cases (Kang et al., 2023; Wang et al., 2024a). ChatUniTest (Chen et al., 2024a) adaptively constructs dependency contexts for the focal method, and employs a generate-validate-fix mechanism to address errors in the generated test cases. Similarly, ChatTester (Yuan et al., 2024) enhances the quality of generated test cases via intent comprehension and iterative correction. While both approaches mitigate the absence of certain dependency information, they fail to account for the influence of project dependencies, such as environment requirements, on the process of test case generation, leaving opportunities for performance optimization. SymPrompt (Ryan et al., 2024), on the other hand, guides LLMs to generate high-coverage test suite code by incorporating symbolic execution-based path information into the prompts. However, SymPrompt does not address the issue of path reachability, which can result in the inclusion of unreachable path information in the prompts, ultimately leading to inaccurate model reasoning. HITS (Wang et al., 2024b) simplifies the analysis of the focal method through program decomposition, achieving high coverage scores by prompting LLMs to generate test cases for sliced code segments. In this work, our proposed framework not only performs an in-depth analysis of the project under test but also utilizes language-specific knowledge to ensure the accuracy and reliability of the unit test cases generated by LLMs.
7. Conclusion

### v2

@@ 4.2.2. The Impact of Different Error-Fixing Phases within CITYWALK
5. Discussion
5.1. Effectiveness of CITYWALK in Bug Detection
Detecting real software bugs is a critical criterion for evaluating the effectiveness of automated unit test generation approaches. To complement our assessment based on correctness and coverage metrics, we employ mutation testing to measure the bug-detection capability of CITYWALK-generated test cases. Prior studies (Jia and Harman, 2011; Fraser and Arcuri, 2011) have validated the utility of mutation testing for both evaluating test quality and guiding the generation of more robust test cases. Mutation testing works by introducing small artificial bugs (i.e., mutants) into the program under test. A test suite is considered effective if it can distinguish the mutated version from the original, i.e., “kill” the mutant by triggering observable failures. In this study, we adopt the updated open-source tool universalmutator (Deb et al., 2024), which supports a broad range of mutation operations, as outlined below:
  • Arithmetic operator mutations: e.g., ++ $\leftrightarrow$ -, $\ast$ $\leftrightarrow$ //
  • Comparison operator mutations: e.g., << $\leftrightarrow$ >>, ==== $\leftrightarrow$ !=!=
  • Logical operator mutations: e.g., && $\leftrightarrow$ $||$
  • Control structure mutations: e.g., removing else, break $\leftrightarrow$ continue
  • Structural mutations: e.g., deleting or commenting out code blocks
  • Literal and constant mutations: e.g., true $\leftrightarrow$ false, 0 $\leftrightarrow$ 1, string substitutions
To facilitate analysis, we select two single-file projects, tinyxml2 and json.cpp, as evaluation subjects. We first perform mutation testing on the test cases generated by CITYWALK, and then calculate the mutation score, which is defined as the ratio of killed mutants to total valid mutants. A higher mutation score indicates stronger bug detection capabilities and higher oracle quality (Xie, 2006). We yield the following key observations according to the evaluation results presented in Table 18:
  1. (1)
    CITYWALK-generated test cases respective achieve a mutation score of 89.34% for json.cpp and 81.12% for tinyxml2. These results indicate that CITYWALK is highly effective in generating test cases that detect a substantial proportion of artificial bugs introduced into the code, validating its capability to uncover potential bugs.
  2. (2)
    The mutation score for json.cpp (created after the GPT-4o knowledge cutoff) outperforms tinyxml2 (which has a risk of potential data leakage) by 8.22%. This provides additional evidence that CITYWALK’s ability to generate high-quality test cases is primarily driven by the designed components as guidance, rather than relying on GPT-4o’s memorization of training data.
  3. (3)
    According to the results from the SBST 2022 tool competition (Schweikl et al., 2022), EvoSuite achieved a mutation score of 34.1% on the competition benchmark. Additionally, the open-source tool universalmutator achieved an average mutation score of 35% on the evaluated C++ projects in the corresponding paper (Deb et al., 2024). In comparison, the mutation scores achieved by CITYWALK are notably higher, further emphasizing the effectiveness of CITYWALK-generated test cases.
Table 18. Mutation Testing for CITYWALK-Generated Test Cases on tinyxml2 and json.cpp
Project # Total Valid Mutants # Killed Mutants Mutation Score
tinyxml2 6272 5088 81.12%
json.cpp 844 754 89.34%
Furthermore, we conduct a case-by-case analysis to assess the usefulness of CITYWALK in discovering real-world bugs. We manually inspect two bugs from tinyxml2 that were identified by CITYWALK-generated test cases. These are real-world issues that have been reported by human developers. As illustrated in Figure 11(a), the CITYWALK-generated test case first creates an XMLDocument object, doc, which then invokes the Value function. This function, in turn, calls XMLNode::Value(). Since the variable _value within XMLNode::Value() may not have been initialized, it triggers an assertion error in the GetStr function. This bug corresponds to Issue #323151515https://github.com/leethomason/tinyxml2/issues/323 in tinyxml2, where a human developer also commits an assertion failure when calling Value. Figure 11(b) shows another bug triggered by CITYWALK, where invalid hexadecimal format string arguments are passed to ToInt64. This bug is associated with Issue #825161616https://github.com/leethomason/tinyxml2/issues/825 in tinyxml2, where a human developer also identifies a similar issue when calling the static member function ToInt64. In summary, the test cases generated by CITYWALK demonstrate the potential to detect real bugs in open-source projects.
@@ 5.1. Effectiveness of CITYWALK in Bug Detection
5.2. Efficiency of CITYWALK
To evaluate the efficiency and practical feasibility of CITYWALK in the context of C++ unit test generation, Table 19 presents the average execution time per focal method and the corresponding token usage for CITYWALK and each baseline approach. The following insights can be drawn from the statistical results:
  1. (1)
    Despite not being the fastest or most cost-efficient approach, CITYWALK maintains acceptable computational overhead. The average execution time per focal method is 34.59 seconds, significantly lower than the slowest baseline (HITS at 184.39 seconds). Since writing unit tests is time-consuming (Fucci et al., 2018), CITYWALK remains a viable option for developers to enhance testing accuracy while ensuring prompt response times. Moreover, although CITYWALK incurs a higher average token usage (5726 tokens per method) compared to most baselines, the incremental cost per method is approximately \$0.03 higher than the best baseline TestPilot, which may be justifiable given CITYWALK’s improved correctness and coverage—especially in scenarios where test reliability is prioritized (e.g., safety-critical systems).
  2. (2)
    Although CITYWALK does not achieve the lowest values in terms of efficiency metrics, it consistently delivers superior test quality, stemming from the designed contextual guidance and post-processing mechanisms. This indicates that CITYWALK strategically trades a moderate increase in computational and monetary cost for substantial gains in test effectiveness, a trade-off often justified in practical software development scenarios.
Table 19. Efficiency Comparison of CITYWALK against the Baselines
Approach CodeGeeX4 DeepSeek-V3 GPT-3.5 GPT-4o ChatTester HITS TestPilot CITYWALK
Avg. Execution Time (s) 15.93 57.02 10.26 18.62 70.30 184.39 24.45 34.59
Avg. Token Usage 2533 3553 1849 2583 $22$ 11736 1436 5726
5.3. Readability and Usability of Test Cases Generated by CITYWALK
The ultimate goal of automated unit test generation is to assist developers in writing test cases. To compare the readability and usability of the test cases generated by CITYWALK with those produced by four LLM-based baselines (i.e., CodeGeeX4, DeepSeek-V3, GPT-3.5, and GPT-4o), we conduct a human evaluation to determine developer preference. We invite five participants, each with over three years of C++ development experience, to perform the assessment. We focus on focal methods for which both CITYWALK and the LLMs generated correct test cases, as recommending test cases that fail to compile or execute is impractical. We further limit our analysis to methods with existing human-written test cases in the original project repositories. Consequently, we randomly select 25 focal methods across five projects from our benchmark. Each participant is required to evaluate 150 test cases, comprising one CITYWALK-generated, four LLM-generated, and one human-written test case for each focal method. Each test case is independently scored across the following four aspects. Scores of each aspect range from 1 (lowest quality) to 3 (highest quality). To avoid bias, participants are not informed about the source of each test case (i.e., whether it is LLM-generated or human-written).
  • Naming Intuitiveness. Clarity and descriptiveness of variable and test method names.
  • Code Layout. Structure, logic, and formatting of the test code.
  • Assertion Quality. Effectiveness and relevance of assertions for validating the focal method.
  • Adoption Efforts. Ease of integrating the test case into real-world usage.
Table 20 summarizes the average scores across all participants. The results show that CITYWALK outperforms all LLM baselines in each of the four aspects. Specifically, CITYWALK surpasses the best baseline by 1.41% in Naming Intuitiveness (compared to DeepSeek-V3), 10.66% in Code Layout (compared to GPT-4o), 18.27% in Assertion Quality (compared to DeepSeek-V3), and 7.14% in Adoption Efforts (compared to DeepSeek-V3 and GPT-4o). Compared with human-written test cases, participants consider CITYWALK-generated test cases with more intuitive naming and more structured code. However, in terms of usability, CITYWALK-generated test cases demonstrate slightly lower assertion quality and comparable adoption effort. This is expected, as generating high-quality assertions remains an open challenge in LLM-based unit test generation. Overall, the findings highlight the superior readability and practical usability of CITYWALK-generated test cases, underscoring its value as a developer-assistive tool.
Table 20. Comparison of CITYWALK against the LLM Baselines in Terms of the Readability and Usability of Generated Test Cases
Approach Readability Usability
Naming Intuitiveness Code Layout Assertion Quality Adoption Efforts
CodeGeeX4 2.70 2.32 1.46 2.47
DeepSeek-V3 2.83 2.39 2.08 2.52
GPT-3.5 2.55 2.26 1.75 2.38
GPT-4o 2.70 2.44 2.06 2.52
CITYWALK 2.87 (1.41% $\uparrow$) 2.70 (10.66% $\uparrow$) 2.46 (18.27% $\uparrow$) 2.70 (7.14% $\uparrow$)
Human-Written 2.64 2.63 2.87 2.75
@@ 5.4. False-Positive Executable Test Cases
Existing LLM-based unit test generation approaches (Yuan et al., 2024; Wang et al., 2024b) primarily rely on compiler error messages to iteratively guide LLMs in fixing test cases that fail during execution. However, this strategy can lead LLMs to align expected values with observed outputs, resulting in test cases that pass without necessarily validating the intended functionality. As illustrated in Figure 12, when provided with the error message, GPT-4o directly uses the expected value ‘‘str’’ as the input for assertions, resulting in a passing test case that is, in reality, a false positive. While such behavior resembles regression testing, a common practice in automated unit test generation, it risks reinforcing incorrect behavior when the underlying implementation contains faults.
Our goal is to highlight this specific risk introduced by compiler message–driven error-fixing strategies in LLM-based frameworks. When LLMs are exposed to runtime or compiler messages that reveal actual outputs, they may inadvertently “learn” to generate assertions that merely reproduce observed behavior rather than verify correctness. Thus, CITYWALK is intentionally designed to avoid relying on compiler feedback for fixing execution failures. Effectively detecting and addressing such false positives remains as our future work.
@@ 5.5. Threats to Validity
In this subsection, we discuss the primary threats to the validity of CITYWALK, as outlined below:
  • External Threat. The primary threats to external validity lie in the diversity of the projects used for evaluation and its generalization to other LLMs. In this paper, we collect 1288 focal methods across ten real-world open-source C++ projects crawled from GitHub, ensuring a degree of diversity and quality in our evaluation. For comparison, we select two open-source code LLMs (CodeGeeX4 and DeepSeek-V3) and two closed-source commercial LLMs (GPT-3.5 and GPT-4o), considering their varied model sizes, architectures, and effectiveness on coding tasks. Future work will involve expanding the benchmark and integrating additional LLMs to better assess the generalizability of CITYWALK.
  • Internal Threat. LLMs exhibit sensitivity to prompt configuration and hyper-parameter settings, particularly the number of task examples and the phrasing of natural language instructions, which can substantially influence performance. To ensure a fair comparison, we use consistent prompts and hyper-parameters across both CITYWALK and all baseline approaches. Additionally, we adopt empirically motivated default configurations rather than fine-tuning prompts or parameters through trial-and-error. We acknowledge that further improvements may be achievable through additional prompt and hyper-parameter tuning. Another potential threat to validity relates to data leakage issue. As GPT-4o is a closed-source model, the exact composition of its training data are not publicly disclosed. Despite this limitation, CITYWALK exhibits a significant improvement in generating high-quality test cases compared to GPT-4o, which utilizes the same underlying architecture. These enhancements suggest that the performance gains achieved by CITYWALK are not merely attributable to the model’s memorization of its training data.
@@ 6.3. LLM-based Automated Unit Test Generation
With the emergence of generative artificial intelligence, researchers have increasingly investigated LLM-driven approaches for automatically generating unit test cases (Kang et al., 2023; Wang et al., 2024a). ChatUniTest (Chen et al., 2024a) adaptively constructs dependency contexts for the focal method, and employs a generate-validate-fix mechanism to address errors in the generated test cases. Similarly, ChatTester (Yuan et al., 2024) enhances the quality of generated test cases via intent comprehension and iterative correction. In contrast to prompt engineering-based approaches that merely append file-level dependencies as additional contextual information, CITYWALK conducts a comprehensive analysis of the project under test to extract project-level dependencies that impact test case generation, such as environment requirements. Moreover, CITYWALK employs RAG techniques to integrate language-specific knowledge from project documentation and source code, thereby enhancing performance. On the other hand, SymPrompt (Ryan et al., 2024) guides LLMs to generate high-coverage test code by incorporating symbolic execution-based path information into the prompts. However, SymPrompt does not address the issue of path reachability, which can result in the inclusion of unreachable path information in the prompts, ultimately leading to inaccurate model reasoning. HITS (Wang et al., 2024b) simplifies the analysis of complex focal methods through program decomposition and achieves high coverage scores by prompting LLMs to generate test cases for sliced code segments. Unlike previous approaches that primarily target interpreted languages such as Java and Python, CITYWALK is specifically designed to address C++-specific challenges. In this work, CITYWALK conducts an empirical analysis to identify common error patterns in LLM-generated C++ unit tests, It then documents these C++-specific failure patterns as insights to guide the post-processing of generated test cases, ensuring their accuracy and reliability.
7. Conclusion and Future Work
@@
Future work will focus on improving LLM-generated unit test quality through two key advancements: (1) developing robust assertion verification techniques to validate functional correctness beyond coverage metrics, ensuring precise alignment with program specifications, and (2) enhancing bug-detection capabilities to identify diverse real-world software faults. These efforts aim to substantially narrow the gap between LLM-generated and human-written tests in terms of reliability and practical utility.
Acknowledgements.
We would like to thank the reviewers for their insightful comments and suggestions. This work was partially supported by the National Key R&D Program of China (Grant No. 2024YFB4505902), the Major Project of ISCAS (Grant No. ISCAS-ZD-202302), the Basic Research Project of ISCAS (Grant No. ISCAS-JCZD-202403), the Youth Innovation Promotion Association of the Chinese Academy of Sciences (Grant Nos. Y2022044 and 2023121).